OpenStudy (anonymous):

Find the direction in which f(x,y) = e^(2x) [cos(3y) - sin(3y)] increases most rapidly at the point (0, pi/2), and find the maximal directional derivative at that point.

OpenStudy (anonymous):

In the direction of the gradient \[gradf(x,y)=\left\{2 e^{2 x} (\cos (3 y)-\sin (3 y)),e^{2 x} (-3 \sin (3 y)-3 \cos (3 y))\right\} \] evaluated at the given point gives \] gradf(0,\pi/2)={2,3} \] The maximum increase is the length of the gradient \[\sqrt{ 2^2 + 3^2}=\sqrt{13} \]